OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. m=5, (4,1) The equation of the line in slope-intercept form is y=?

OpenStudy (anonymous):

y-1=5(x-4), so, y=5x-19

OpenStudy (anonymous):

y-1= 5(x-4) y=5x-19